# On solutions of reduced biquaternion equality constrained least squares problem and their relative forward error bound 22footnotemark: 2

### v1

@@ Abstract
This study focuses on addressing the challenge of solving the reduced biquaternion equality constrained least squares (RBLSE) problem. We develop algebraic techniques to derive both complex and real solutions for the RBLSE problem by utilizing the complex and real forms of reduced biquaternion matrices. Additionally, we conduct a perturbation analysis for the RBLSE problem and establish an upper bound for the relative forward error of these solutions. Numerical examples are presented to illustrate the effectiveness of the proposed approaches and to verify the accuracy of the established upper bound for the relative forward errors.
@@ 1 Introduction
Reduced biquaternions, first introduced by Segre in $1892$, have become fundamental in various domains such as signal processing, image analysis, and neural networks [4, 5, 6, 7]. Unlike standard quaternions, reduced biquaternions, often referred to as commutative quaternions, possess the unique property of commutative multiplication. This property not only simplifies mathematical operations but also significantly boosts computational efficiency, as demonstrated in [9, 10].
@@
In many practical applications, determining the solution to linear system, typically expressed as $AX\approx B$, is a common challenge. The least squares method is a well-known approach for addressing this problem. Prior research, such as the work in [12], has explored least squares solutions for matrix equations like $AX=B$ and $AXC=B$ over commutative quaternions. The equality constrained least squares problem is a further study of the least squares problem, and in [13], the authors discussed solution techniques for computing the reduced biquaternion solution of the reduced biquaternion equality constrained least squares (RBLSE) problem. In this manuscript, we advance the research by developing methods to obtain both complex and real solutions for the RBLSE problem. Additionally, we establish an upper bound for the relative forward error associated with these solutions, ensuring the accuracy and reliability of our method in addressing RBLSE problems.
The organization of the manuscript is as follows: Section 2 covers the notation and preliminary concepts. Section 3 outlines the method for finding complex solution of the RBLSE problem, while Section 4 focuses on the technique for finding real solutions to the same problem. Lastly, Section 5 presents the numerical validation of the proposed methods.
@@ 2.1 Notation
In this paper, we adopt the following notations: $\mathbb{Q}_{\mathbb{R}}^{m\times n}$, ${\mathbb{C}}^{m\times n}$, and ${\mathbb{R}}^{m\times n}$ represent the sets of all $m\times n$ reduced biquaternion, complex, and real matrices. The notations $\left\lVert\cdot\right\rVert_{2}$ and $\left\lVert\cdot\right\rVert_{F}$ refer to the $2$-norm and Frobenius norm, respectively. For matrices $A\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$ and $B\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $\left[A,B\right]$ represents their concatenation, i.e., $\begin{bmatrix}A&B\end{bmatrix}\in\mathbb{Q}_{\mathbb{R}}^{m\times(n+d)}$. For a matrix $A\in{\mathbb{C}}^{m\times n}$, the symbols $A^{T}$, $A^{H}$, and $A^{\dagger}$ correspond to the transpose, conjugate transpose, and Moore-Penrose inverse of $A$, respectively. Additionally, $A^{-1}$ indicates the inverse of $A$ if it exists.
The Matlab functions $randn(m,n)$ and $rand(m,n)$ produce $m\times n$ matrices filled with normally distributed and uniformly distributed random values, respectively. The abbreviations RBLSE (reduced biquaternion equality constrained least squares) and LSE (equality constrained least squares) will be used throughout this paper.
@@ 2.2 Preliminaries
$M^{R}=\begin{bmatrix}M_{0}&-M_{1}&M_{2}&-M_{3}\\ M_{1}&M_{0}&M_{3}&M_{2}\\ M_{2}&-M_{3}&M_{0}&-M_{1}\\ M_{3}&M_{2}&M_{1}&M_{0}\end{bmatrix},\;\;M^{C}=\begin{bmatrix}N_{1}&N_{2}\\ N_{2}&N_{1}\end{bmatrix}.$ (2)
@@
$Q_{m}=\begin{bmatrix}0&-I_{m}&0&0\\ I_{m}&0&0&0\\ 0&0&0&-I_{m}\\ 0&0&I_{m}&0\end{bmatrix},\;R_{m}=\begin{bmatrix}0&0&I_{m}&0\\ 0&0&0&I_{m}\\ I_{m}&0&0&0\\ 0&I_{m}&0&0\end{bmatrix},\mbox{and}\;S_{m}=\begin{bmatrix}0&0&0&-I_{m}\\ 0&0&I_{m}&0\\ 0&-I_{m}&0&0\\ I_{m}&0&0&0\end{bmatrix}.$ (4)
We have
@@
Let $M_{c}^{C}$ represent the first block column of the matrix $M^{C}$. More precisely, $M_{c}^{C}$ is defined as: $M^{C}_{c}=\left[M_{1}^{T},M_{2}^{T}\right]^{T}.$ The matrix $M^{C}$ can be represented using $M^{C}_{c}$ as:
@@
We have
@@ 3 An algebraic technique for complex solution of RBLSE problem
$A$ $=$ $M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\;\;\;\;B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (9)
$C$ $=$ $R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\;\;\;\;D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (10)
We will limit our discussion to the scenario where $m\geq n+d$, and both matrices $C$ and $C^{C}_{c}$ have full row rank. The RBLSE problem can then be stated as follows:
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;CX=D.$ (11)
@@
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{C}_{c}X=D^{C}_{c}.$ (12)
Let the QR factorization of $\left(C^{C}_{c}\right)^{H}$ be as follows:
@@
where $\widetilde{Q}\in{\mathbb{C}}^{n\times n}$ is a unitary matrix and $\widetilde{R}\in{\mathbb{C}}^{2p\times 2p}$ is a nonsingular upper triangular matrix. Now, partition $A^{C}_{c}\widetilde{Q}$ as:
@@
Theorem 3.1.
Consider the RBLSE problem outlined in (11) and the complex LSE problem in (12), with notations specified in (13) and (14). For a matrix $X\in{\mathbb{C}}^{n\times d}$, $X$ is a complex solution of the RBLSE problem (11) if and only if $X$ solves the complex LSE problem (12). In this scenario, the unique solution $X$ with minimum norm can be expressed as:
$X=\widetilde{Q}\begin{bmatrix}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}\\ \widetilde{P}_{2}^{\dagger}\left(B^{C}_{c}-\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}\right)\end{bmatrix}.$ (15)
@@
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\textrm{min},\;\;C^{C}_{c}X=D^{C}_{c}.$ (16)
@@
To find the expression for $X$, we solve the complex LSE problem (12). Set
$\widetilde{Q}^{H}X=\begin{bmatrix}Y\\ Z\end{bmatrix},$
where $Y\in{\mathbb{C}}^{2p\times d}$ and $Z\in{\mathbb{C}}^{(n-2p)\times d}$. Equation 12 can be rewritten as
$\min_{X}\left\lVert A^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X-B^{C}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X=D^{C}_{c}.$ (19)
Utilizing (13), we have
$C^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X=D^{C}_{c}\implies\begin{bmatrix}\widetilde{R}^{H}&0\end{bmatrix}\begin{bmatrix}Y\\ Z\end{bmatrix}=D^{C}_{c}.$
Since $\widetilde{R}^{H}$ is a nonsingular matrix, we get $Y=\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$. Using (14), Equation 19 takes the form
$\min_{Z}\left\lVert\widetilde{P}_{2}Z-\left(B^{C}_{c}-\widetilde{P}_{1}Y\right)\right\rVert_{F}.$
The minimum norm solution of the above least squares problem is $Z=\widetilde{P}_{2}^{\dagger}\left(B^{C}_{c}-\widetilde{P}_{1}Y\right)$. Thus, we can derive the desired expression for $X$. ∎
@@
$\widehat{A}=A+\Delta A,\;\widehat{B}=B+\Delta B,\;\widehat{C}=C+\Delta C,\;\textrm{and}\;\widehat{D}=D+\Delta D,$ (20)
where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively. We assume that the perturbation $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to guarantee that the perturbed matrices $\widehat{C}$ and $\widehat{C}^{C}_{c}$ retain full row rank. These perturbations are measured normwise by the smallest $\epsilon$ for which
$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\left\lVert\Delta C\right\rVert_{F}\leq\epsilon\left\lVert C\right\rVert_{F},\;\textrm{and}\;\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (21)
@@
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\;\;\textrm{subject to}\;\;\widehat{C}X=\widehat{D},$ (22)
@@
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{C}_{c}X-\left(\widehat{B}\right)^{C}_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(\widehat{C}\right)^{C}_{c}X=\left(\widehat{D}\right)^{C}_{c}.$ (24)

### v2

# On the specific solutions of reduced biquaternion equality constrained least squares problem and their relative forward error bound 22footnotemark: 2

@@ Abstract
This study focuses on addressing the challenge of solving the reduced biquaternion equality constrained least squares (RBLSE) problem. We develop algebraic techniques to derive real and complex solutions for the RBLSE problem by utilizing the real and complex forms of reduced biquaternion matrices. Furthermore, we propose algorithms and provide a detailed analysis of their computational complexity for finding special solutions to the RBLSE problem. A perturbation analysis is conducted, establishing an upper bound for the relative forward error of these solutions. This analysis ensures the accuracy and stability of the solutions in the presence of data perturbations, which is crucial for practical applications where errors arising from input inaccuracies can cause deviations between computed and true solutions. Numerical examples are presented to validate the proposed algorithms, demonstrate their effectiveness, and verify the accuracy of the established upper bound for the relative forward errors. These findings lay the groundwork for exploring applications in 3D and 4D algebra such as robotics, signal, and image processing, expanding their impact on practical and emerging domains.
@@ 1 Introduction
Reduced biquaternions, first introduced by Segre in $1892$, have become fundamental in various domains such as signal processing, image processing, and neural networks [5, 6, 7, 9]. Unlike standard quaternions, reduced biquaternions, often referred to as commutative quaternions, possess the unique property of commutative multiplication. This property not only simplifies mathematical operations but also significantly boosts computational efficiency, as demonstrated in [12, 13].
Reduced biquaternions also exhibit significant differences in their algebraic structure compared to traditional quaternions. Two notable elements, $e_{1}$ and $e_{2}$, are defined as:
$e_{1}=\frac{1+\textit{{j}}}{2},\quad e_{2}=\frac{1-\textit{{j}}}{2}.$
These elements satisfy $e_{1}e_{2}=0$, $e_{1}^{2}=e_{1}$, $e_{2}^{2}=e_{2}$, and are idempotent as well as divisors of zero. Any reduced biquaternion element of the form $c_{1}e_{1}$ or $c_{2}e_{2}$ (where $c_{1}$ and $c_{2}$ are complex numbers) also acts as a divisor of zero, highlighting the non-division algebra property of reduced biquaternions. Consequently, $\mathbb{Q}_{\mathbb{R}}$, the reduced biquaternion algebra, does not form a complete division algebra.
The concept of units in $\mathbb{Q}_{\mathbb{R}}$ is also unique. An element $r\in\mathbb{Q}_{\mathbb{R}}$ is defined as a unit if there exists another element $s\in\mathbb{Q}_{\mathbb{R}}$ such that $rs=sr=1$. However, not every nonzero element in $\mathbb{Q}_{\mathbb{R}}$ is a unit, unlike in ${\mathbb{R}}$, ${\mathbb{C}}$ or in quaternions. For instance, $e_{1}$ or $e_{2}$ is not a unit in $\mathbb{Q}_{\mathbb{R}}$. These differences underline the distinct algebraic framework of reduced biquaternions and motivate their applicability to specific problems in mathematical and physical sciences.
While reduced biquaternions have several advantages, their non-division algebra nature has negligible impact on practical applications. For instance, in [6, 9, 12, 13], the authors demonstrated the superior performance of reduced biquaternions over traditional quaternions in signal and image processing. This makes reduced biquaternions particularly advantageous for real-time applications in these domains. Moreover, their applicability extends beyond numerical computations to areas such as theoretical physics, including their connection to Maxwell’s equations [1, 10].
@@
In many practical applications, determining the solution to linear system, typically expressed as $AX\approx B$, is a common challenge. The least squares method is a well-known approach for addressing this problem. Prior research, such as the work in [16], has explored least squares solutions for matrix equations like $AX=B$ and $AXC=B$ over commutative quaternions. The equality constrained least squares problem is a further study of the least squares problem, and in [17], the authors discussed solution techniques for computing the reduced biquaternion solution of the reduced biquaternion equality constrained least squares (RBLSE) problem. Recent work has extended these studies to quaternions and split quaternions, proposing algorithms for solving split quaternion and quaternion equality constrained least squares problems [14, 20].
Despite these advancements, existing studies primarily focus on computational techniques for finding general solutions to LSE problems over quaternions, split quaternions, and reduced biquaternions. However, specialized solutions to the RBLSE problem—such as real and complex solutions—have not been explored. Furthermore, the perturbation analysis of these solutions have not been addressed.
In numerical analysis, the concept of relative forward error is central to evaluating the accuracy and stability of computed solutions, particularly for the RBLSE problem. This metric is crucial in the presence of data perturbations, where errors from machine precision, floating-point arithmetic, or input inaccuracies can lead to deviations between computed and true solutions. The relative forward error quantifies these deviations, providing insights into the sensitivity of solutions to data perturbations and enabling the assessment of solution accuracy and reliability. Understanding these errors is essential in practical applications, such as robotics, image processing, and control systems, where reduced biquaternion algebra plays a key role.
To address these gaps, this study aims to:
  1. 1.
    Develop algebraic techniques to obtain real and complex solutions of the RBLSE problem.
  2. 2.
    Establish upper bounds for the relative forward error associated with these solutions to ensure accuracy and stability.
  3. 3.
    Propose efficient algorithms for solving the RBLSE problem and analyze their computational complexity.
  4. 4.
    Conduct numerical experiments to validate the proposed methods, assess their computational efficiency, and verify the accuracy of the established upper bounds.
Among the proposed solutions, the computational complexity analysis reveals that finding the real solution to the RBLSE problem requires less computational time than obtaining the complex solution. Numerical experiments validate the computational efficiency of the algorithms, while the established upper bounds for the relative forward error provide insights into solution sensitivity and reliability under perturbations.
These contributions are vital for advancing the numerical analysis of reduced biquaternions and ensuring the robustness of solutions in applications such as signal processing, color image processing, and robotics. By addressing both computational techniques and perturbation analysis, this study lays the groundwork for further exploration of reduced biquaternion algebra in practical and emerging domains.
The organization of the manuscript is as follows: Section 2 covers the notation and preliminary concepts. Section 3 outlines the method for finding the real solution of the RBLSE problem, while Section 4 focuses on the technique for finding complex solutions to the same problem. Lastly, Section 5 presents the numerical validation of the proposed methods.
@@ 2.1 Notation
In this paper, we adopt the following notations: $\mathbb{Q}_{\mathbb{R}}^{m\times n}$, ${\mathbb{C}}^{m\times n}$, and ${\mathbb{R}}^{m\times n}$ represent the sets of all $m\times n$ reduced biquaternion, complex, and real matrices. The notation $\left\lVert\cdot\right\rVert_{F}$ represents the Frobenius norm, which applies to matrices over real numbers, complex numbers, or reduced biquaternions, while $\left\lVert\cdot\right\rVert_{2}$ denotes the $2$-norm, specifically for real or complex matrices. For matrices $A\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$ and $B\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $\left[A,B\right]$ represents their concatenation, i.e., $\begin{bmatrix}A&B\end{bmatrix}\in\mathbb{Q}_{\mathbb{R}}^{m\times(n+d)}$. For a matrix $A\in{\mathbb{C}}^{m\times n}$, the symbols $A^{T}$, $A^{H}$, and $A^{\dagger}$ correspond to the transpose, conjugate transpose, and Moore-Penrose inverse of $A$, respectively. Additionally, $A^{-1}$ indicates the inverse of $A$ if it exists.
The Matlab function $rand(m,n)$ generates an $m\times n$ matrix filled with uniformly distributed random values. The abbreviations RBLSE (reduced biquaternion equality constrained least squares) and LSE (equality constrained least squares) will be used throughout this paper.
@@ 2.2 Preliminaries
$M^{R}=\begin{bmatrix}M_{0}&-M_{1}&M_{2}&-M_{3}\\ M_{1}&M_{0}&M_{3}&M_{2}\\ M_{2}&-M_{3}&M_{0}&-M_{1}\\ M_{3}&M_{2}&M_{1}&M_{0}\end{bmatrix},\quad M^{C}=\begin{bmatrix}N_{1}&N_{2}\\ N_{2}&N_{1}\end{bmatrix}.$ (2)
@@
$Q_{m}=\begin{bmatrix}0&-I_{m}&0&0\\ I_{m}&0&0&0\\ 0&0&0&-I_{m}\\ 0&0&I_{m}&0\end{bmatrix},\quad R_{m}=\begin{bmatrix}0&0&I_{m}&0\\ 0&0&0&I_{m}\\ I_{m}&0&0&0\\ 0&I_{m}&0&0\end{bmatrix},\quad S_{m}=\begin{bmatrix}0&0&0&-I_{m}\\ 0&0&I_{m}&0\\ 0&-I_{m}&0&0\\ I_{m}&0&0&0\end{bmatrix}.$ (4)
The Frobenius norm, as described in [16, 17], is defined using the real representation of reduced biquaternion matrices. This definition aligns with the definition provided in (1), ensuring consistency. Specifically, the Frobenius norm of $M$ satisfies:
@@
Let $M_{c}^{C}$ represent the first block column of the matrix $M^{C}$. More precisely, $M_{c}^{C}$ is defined as: $M^{C}_{c}=\left[N_{1}^{T},N_{2}^{T}\right]^{T}.$ The matrix $M^{C}$ can be represented using $M^{C}_{c}$ as:
@@
The Frobenius norm, as described in [16, 17], is also expressed using the complex representation of reduced biquaternion matrices. This definition is consistent with the previously stated definition (1). Specifically, the Frobenius norm of $M$ satisfies:
@@ 4. $(d)$
In the context of solving constrained least squares problems, the following result provides the closed-form solution for the complex LSE problem. This theorem, based on the QR factorization technique, is foundational for deriving efficient algorithms for more generalized problems discussed later in this paper. The result is formally stated as follows:
Lemma 2.4.
[8] Consider the complex LSE problem:
$\min_{X}\|AX-B\|_{F}\quad\text{subject to}\quad CX=D,$ (9)
where $A\in\mathbb{C}^{m\times n}$, $B\in\mathbb{C}^{m\times d}$, $C\in\mathbb{C}^{p\times n}$, and $D\in\mathbb{C}^{p\times d}$. Assume that the matrix $C$ has full row rank. Let the QR factorization of $C^{H}$ be given by:
$C^{H}=Q\begin{bmatrix}R\\ 0\end{bmatrix},$ (10)
where $Q\in\mathbb{C}^{n\times n}$ is a unitary matrix and $R\in\mathbb{C}^{p\times p}$ is a nonsingular upper triangular matrix. Further, partition the product $AQ$ as:
$AQ=\left[P_{1},P_{2}\right],$ (11)
where $P_{1}\in\mathbb{C}^{m\times p}$ and $P_{2}\in\mathbb{C}^{m\times(n-p)}$.
The unique solution $X\in\mathbb{C}^{n\times d}$ with minimum norm for the complex LSE problem (9) is given by:
$X=Q\begin{bmatrix}\left(R^{H}\right)^{-1}D\\ P_{2}^{\dagger}\left(B-P_{1}\left(R^{H}\right)^{-1}D\right)\end{bmatrix}.$ (12)
Proof.
Equation (9) can be rewritten as
$\min_{X}\left\lVert AQQ^{H}X-B\right\rVert_{F}\quad\textrm{subject to}\quad CQQ^{H}X=D.$ (13)
Set
$Q^{H}X=\begin{bmatrix}Y\\ Z\end{bmatrix},$
where $Y\in\mathbb{C}^{p\times d}$ and $Z\in\mathbb{C}^{(n-p)\times d}$. Utilizing (10), we have
$CQQ^{H}X=D\implies\begin{bmatrix}R^{H}&0\end{bmatrix}\begin{bmatrix}Y\\ Z\end{bmatrix}=D.$
Since $R^{H}$ is a nonsingular matrix, we get $Y=\left(R^{H}\right)^{-1}D$. Using (11), equation (13) takes the form
$\min_{Z}\left\lVert P_{2}Z-\left(B-P_{1}Y\right)\right\rVert_{F}.$
The minimum norm solution of the above least squares problem is $Z=P_{2}^{\dagger}\left(B-P_{1}Y\right)$.
Finally, substituting $Y$ and $Z$ back into $Q^{H}X=\begin{bmatrix}Y\\ Z\end{bmatrix}$, we can derive the desired expression for $X$. This completes the proof. ∎
3 An algebraic technique for real solution of RBLSE problem
$A$ $=$ $A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\quad B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (14)
$C$ $=$ $C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\quad D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (15)
We will limit our discussion to the scenario where $m\geq n+d$, and matrix $C^{R}_{c}$ has full row rank. The RBLSE problem can then be stated as follows:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\quad\textrm{subject to}\quad CX=D.$ (16)
Consider a real LSE problem
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}\quad\textrm{subject to}\quad C^{R}_{c}X=D^{R}_{c}.$ (17)
Theorem 3.1.
Consider the RBLSE problem outlined in (16) and the real LSE problem in (17). Let the QR factorization of $\left(C^{R}_{c}\right)^{T}$ be given by:
where $\bar{Q}\in\mathbb{R}^{n\times n}$ is an orthonormal matrix and $\bar{R}\in\mathbb{R}^{4p\times 4p}$ is a nonsingular upper triangular matrix. Further, partition the product $A^{R}_{c}\bar{Q}$ as:
$A^{R}_{c}\bar{Q}=\left[\bar{P}_{1},\bar{P}_{2}\right],$ (19)
Let $X\in\mathbb{R}^{n\times d}$. Then, $X$ is a real solution of the RBLSE problem (16) if and only if $X$ solves the real LSE problem (17). In this scenario, the unique solution $X$ with minimum norm can be expressed as:
Proof.
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\textrm{min},\quad C^{R}_{c}X=D^{R}_{c}.$ (21)
Using equations (2), (3), (5), along with Lemmas 2.1 and 2.3, we obtain
$=$ $\frac{1}{2}\left\lVert A^{R}X^{R}-B^{R}\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert\left(AX-B\right)^{R}\right\rVert_{F}$
$=$ $\left\lVert AX-B\right\rVert_{F}.$
From (21), we obtain
and
$C^{R}X^{R}$ $=$ $D^{R},$
$(CX)^{R}$ $=$ $D^{R},$
$CX$ $=$ $D.$ (23)
By combining (22) and (23), we conclude that $X\in{\mathbb{R}}^{n\times d}$ is a real solution to the RBLSE problem (16), and vice versa.
To determine the expression for $X$, we solve the real LSE problem given in (17). Using Lemma 2.4, we derive the explicit expression for $X$. ∎
$\widehat{A}=A+\Delta A,\quad\widehat{B}=B+\Delta B,\quad\widehat{C}=C+\Delta C,\quad\textrm{and}\quad\widehat{D}=D+\Delta D,$ (24)
where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively. We assume that the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to ensure that the perturbed matrix $\widehat{C}^{R}_{c}$ retains full row rank. These perturbations are measured normwise by the smallest $\epsilon$ for which
$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\quad\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\quad\left\lVert\Delta C\right\rVert_{F}\leq\epsilon\left\lVert C\right\rVert_{F},\quad\textrm{and}\quad\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (25)
Let $\widehat{X}_{RL}$ be the real solution to the perturbed RBLSE problem
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\quad\textrm{subject to}\quad\widehat{C}X=\widehat{D},$ (26)
and let $\Delta X_{RL}=\widehat{X}_{RL}-X_{RL}$.
Theorem 3.2.
where
Proof.
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{R}_{c}X-\left(\widehat{B}\right)^{R}_{c}\right\rVert_{F}\quad\textrm{subject to}\quad\left(\widehat{C}\right)^{R}_{c}X=\left(\widehat{D}\right)^{R}_{c}.$ (28)
$\left(\widehat{A}\right)^{R}_{c}=A^{R}_{c}+\left(\Delta A\right)^{R}_{c},\quad\left(\widehat{B}\right)^{R}_{c}=B^{R}_{c}+\left(\Delta B\right)^{R}_{c},\quad\left(\widehat{C}\right)^{R}_{c}=C^{R}_{c}+\left(\Delta C\right)^{R}_{c},\quad\left(\widehat{D}\right)^{R}_{c}=D^{R}_{c}+\left(\Delta D\right)^{R}_{c}.$
Thus, the perturbed real LSE problem (28) can be rewritten as:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(A^{R}_{c}+\left(\Delta A\right)^{R}_{c}\right)X-\left(B^{R}_{c}+\left(\Delta B\right)^{R}_{c}\right)\right\rVert_{F}\quad\textrm{subject to}\quad\left(C^{R}_{c}+\left(\Delta C\right)^{R}_{c}\right)X=\left(D^{R}_{c}+\left(\Delta D\right)^{R}_{c}\right).$ (29)
$\left\lVert\left(\Delta A\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{R}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta B\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{R}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta C\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{R}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta D\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{R}_{c}\right\rVert_{F}.$ (30)
With the perturbed problem (29) and the bounds in (30), and using Theorem 3.1, the sensitivity analysis of the real solution to the RBLSE problem (16) reduces to evaluating the sensitivity of the solution to the real LSE problem (17). Consequently, the upper bound $U_{RL}$ for the relative forward error of the real solution to the RBLSE problem can be obtained from [2, Equation $4.11$]. ∎
@@ 4 An algebraic technique for complex solution of RBLSE problem
$A$ $=$ $M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\quad B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (31)
$C$ $=$ $R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\quad D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (32)
We will limit our discussion to the scenario where $m\geq n+d$, and matrix $C^{C}_{c}$ has full row rank. The RBLSE problem can then be stated as follows:
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\quad\textrm{subject to}\quad CX=D.$ (33)
@@
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}\quad\textrm{subject to}\quad C^{C}_{c}X=D^{C}_{c}.$ (34)
Theorem 4.1.
Consider the RBLSE problem outlined in (33) and the complex LSE problem in (34). Let the QR factorization of $(C^{C}_{c})^{H}$ be given by:
@@
where $\widetilde{Q}\in\mathbb{C}^{n\times n}$ is a unitary matrix and $\widetilde{R}\in\mathbb{C}^{2p\times 2p}$ is a nonsingular upper triangular matrix. Further, partition the product $A^{C}_{c}\widetilde{Q}$ as:
@@
Let $X\in\mathbb{C}^{n\times d}$. Then, $X$ is a complex solution of the RBLSE problem (33) if and only if $X$ solves the complex LSE problem (34). In this scenario, the unique solution $X$ with minimum norm can be expressed as:
$X=\widetilde{Q}\begin{bmatrix}(\widetilde{R}^{H})^{-1}D^{C}_{c}\\ \widetilde{P}_{2}^{\dagger}\big{(}B^{C}_{c}-\widetilde{P}_{1}(\widetilde{R}^{H})^{-1}D^{C}_{c}\big{)}\end{bmatrix}.$ (37)
@@
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\textrm{min},\quad C^{C}_{c}X=D^{C}_{c}.$ (38)
@@
To find the expression for $X$, we solve the complex LSE problem (34). Using Lemma 2.4, we derive the explicit expression for $X$. ∎
@@
$\widehat{A}=A+\Delta A,\quad\widehat{B}=B+\Delta B,\quad\widehat{C}=C+\Delta C,\quad\textrm{and}\quad\widehat{D}=D+\Delta D,$ (41)
where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively. We assume that the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to guarantee that the perturbed matrix $\widehat{C}^{C}_{c}$ retains full row rank. These perturbations are measured normwise by the smallest $\epsilon$ for which
$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\quad\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\quad\left\lVert\Delta C\right\rVert_{F}\leq\epsilon\left\lVert C\right\rVert_{F},\quad\textrm{and}\quad\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (42)
@@
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\quad\textrm{subject to}\quad\widehat{C}X=\widehat{D},$ (43)
@@
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{C}_{c}X-\left(\widehat{B}\right)^{C}_{c}\right\rVert_{F}\quad\textrm{subject to}\quad\left(\widehat{C}\right)^{C}_{c}X=\left(\widehat{D}\right)^{C}_{c}.$ (45)